# Optimized Many-Hypercube Codes toward Lower Logical Error Rates and Earlier Realization

### v1

@@ Abstract
Many-hypercube codes [H. Goto, Sci. Adv. 10, eadp6388 (2024)], concatenated [[n,n2,2]]{[[n,n-2,2]]} quantum error-detecting codes ($n$ is even), have recently been proposed as high-rate quantum codes suitable for fault-tolerant quantum computing. However, the original many-hypercube codes with n=6{n=6} have large code block sizes at high concatenation levels (216 and 1296 physical qubits per block at levels 3 and 4, respectively), making not only experimental realization difficult but also logical error rates high. Toward earlier experimental realization and lower logical error rates, here we investigate smaller many-hypercube codes obtained by concatenating $[[6,4,2]]$ and/or $[[4,2,2]]$ codes, where, e.g., $D_{6,4,4}$ denotes the many-hypercube code using $[[6,4,2]]$ at level 1 and $[[4,2,2]]$ at levels 2 and 3. As a result, we found a surprising fact: $D_{6,4,4}$ ($D_{6,6,4,4}$) can achieve lower block error rates than $D_{4,4,4}$ ($D_{4,4,4,4}$), despite its higher encoding rate. Focusing on level 3, we also developed efficient fault-tolerant encoders realizing about 60% overhead reduction while maintaining or even improving the performance, compared to the original design. Using them, we numerically confirmed that $D_{6,4,4}$ also achieves the best performance for logical controlled-NOT gates in a circuit-level noise model. These results will be useful for early experimental realization of fault-tolerant quantum computing with high-rate quantum codes.
@@ I Introduction
Topological codes [24], in particular, the surface code [25, 26] have become more popular since it was shown that the surface code can achieve comparable thresholds only using local operations [27, 28, 29, 30, 31]. Moreover, a compact version called the rotated surface code and a simple two-qubit gate approach called lattice surgery were also proposed [32]. The surface-code approach to FTQC is particularly suitable for superconducting-circuit implementations, leading to recent experimental demonstrations of the surface code with superconducting circuits [33, 34, 35, 36].
@@
The (rotated) surface code encodes a single logical qubit into $d^{2}$ physical qubits, where $d$ is the code distance. Thus, the encoding rate $R=1/d^{2}$ becomes low for large code block sizes, resulting in large resource overheads. To address this overhead issue, quantum low-density parity-check (qLDPC) codes [43, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], which can achieve a constant encoding rate even for large code distances, have recently attracted much attention [48, 49, 50, 51, 52, 53]. This expectation is enhanced by recent experimental advances, in particular, in ion-trap [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] and neutral-atom [68, 69, 70, 71, 72, 73, 74, 75, 76, 77] platforms, because long-range interactions beyond nearest-neighbor ones necessary for high-rate codes such as qLDPC codes [78, 79] are experimentally feasible by moving qubits in these platforms. However, parallel execution of addressable logical gates with low overheads is still challenging for qLDPC codes, the study on which is ongoing [80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102].
@@
Alternative high-rate codes, namely, non-LDPC high-rate quantum codes have also been proposed toward low-overhead FTQC [103, 104, 105, 106, 107, 108, 109, 110]. Among them, concatenated high-rate quantum codes are particularly promising in the sense that fault-tolerant logical-state preparation can be achieved by level-by-level encoding [16, 22, 64, 103, 104, 105]. In this work, we focus on concatenated [[n,n2,2]]{[[n,n-2,2]]} quantum error-detecting codes ($n$ is even), which are called many-hypercube (MHC) codes because the code structure can be described by using many $L$-dimensional hypercubes ($L$ is the concatenation level) corresponding to logical qubits (logical Pauli operators) [104, 107, 109]. Note that the error-detecting codes, which are also called iceberg codes [60], are the simplest, highest-rate Calderbank-Shor-Steane (CSS) codes [3, 4, 5]. In the original work [104], concatenated $[[6,4,2]]$ codes were adopted because of the good balance between the high rate and code block size of $[[6,4,2]]$. Recently, smaller MHC codes with $[[4,2,2]]$ have been studied [107, 109]. Remarkably, the smallest level-2 MHC code $4^{2},2^{2},2^{2}]]$ has been realized experimentally using neutral atoms [77]. Smaller block sizes will lead to not only earlier experimental realization, but also lower logical error rates, though the encoding rates become lower. For this reason, in this work we investigate the MHC codes with $[[6,4,2]]$ and/or $[[4,2,2]]$, where the MHC code with $n_{L},n_{L}-2,2]]$} at level $L$ is denoted by $D_{n_{1},n_{2},\ldots}$. As a result, we found that $D_{6,4,4}$ ($D_{6,6,4,4}$) can achieve lower block error rates than $D_{4,4,4}$ ($D_{4,4,4,4}$), despite its higher encoding rate. In addition, it turned out that $D_{4,4,6,6}$ is the worst at level 4. This is a notable result, because in a previous study [107], it was assumed that we should use smaller codes at lower levels, and consequently $D_{4,4,6,6}$ was adopted in the previous study. We also develop efficient fault-tolerant encoders for level-3 MHC codes, which can reduce overheads by about 60% compared to the original design [104]. Using them, we confirmed that $D_{6,4,4}$ can also achieve the highest performance for logical CNOT gates at level 3.
@@
This paper is organized as follows. In Sec. II, we explain the definitions of the MHC codes in detail using figures visualizing their code structures. In Sec. III, we show numerical results for block error rates using a bit-flip (depolarizing) error model. In Sec. IV, we present our new encoders for level-3 MHC codes comparing them with the original ones. In Sec. V, we show numerical results for logical-CNOT error rates at level 3 using the new and original encoders. Finally, we summarize the results and discuss the outlook in Sec. VI.
@@ II.1 $D_{4}$ and $D_{6}$
Similarly to the case of $D_{4}$, these can be visualized as shown in Fig. 1(b). There are different definitions of [[n,n2,2]][[n,n-2,2]] codes [60], we adopt the above definitions because of the simple visualization [104].
@@ II.4 Level-4 MHC codes
The level-4 MHC codes $D_{n_{1},n_{2},n_{3},4}$ and $D_{n_{1},n_{2},n_{3},6}$ are defined by similar equations to those for $D_{n_{1},n_{2},4}$ and $D_{n_{1},n_{2},6}$, respectively, in the last sectoin. By introducing the fourth virtual axis denoted by $t$ and stacking level-3 $D_{n_{1},n_{2},n_{3}}$ blocks along the $t$ axis, each logical Pauli operator of the level-4 MHC codes corresponds to a four-dimensional hypercube, where $\mathrm{Q}_{i,j,k}$ of the $l$th $D_{n_{1},n_{2},n_{3}}$ block is placed at $(x,y,z,t)=(i,j,k,l)$. This is the origin of the name of MHC codes.
@@ III Performance comparison of MHC codes
Thus, the initial state can be prepared using the encoders in Figs. 5(a) and 5(b), respectively, where we assume no errors on all the operations. Next, we induce bit-flip errors on the physical qubits, then measure all the physical qubits in the $Z$ basis, and finally decode the measurement results using the level-by-level minimum-distance decoder proposed in the original work [104, 111]. Unless all the logical values obtained by the decoding are 0, we consider that the decoding fails, leading to the block error probability. In this work, we used a Python package called Stim [115] for quantum circuit simulations.
@@
We also evaluated the performance of level-3 and level-4 MHC codes at $p_{\mathrm{flip}}=1\%$} and 2%, respectively, the results of which are shown in Figs. 5(b) and 5(c), respectively. As in the case of level 2, our result clearly suggests that we should use larger-size codes for lower levels for MHC codes, contrary to the previous assumption [107]. More importantly, Figs. 5(b) and 5(c) provide another counterintuitive result that the highest-performance MHC code is not the smallest-size one. The best MHC code is $D_{6,4,4}$ at level 3 and $D_{6,6,4,4}$ at level 4. (The level-4 MHC code chosen in a previous work [107], $D_{4,4,6,6}$, is the worst.) This is the main result of this work.
@@ IV Proposed encoders of level-3 MHC codes
$=n_{2}n_{1}+n_{2}+3n_{1}/2-2$ (4)
for the proposed encoder. (Note that the latter formula holds only for $n_{1},n_{2}=4,6$}.) Thus, the proposed encoder can reduce the overhead by $7-n_{1}/2$.
@@
Figure 7: Direct measurements of level-2 logical $Z$ operators. Here we show only the cases of $D_{i,4}$ (i=4,6{i=4,6}), but the cases of $D_{i,6}$ are similar. The squares are physical ancilla qubits used for the measurements, which are initially set in $|0\rangle$. (a,b) Gate sequence for the logical $Z$ measurements used in the $D_{4,4}$ and $D_{6,4}$ encoders, respectively. Note that the bottom two level-1 blocks correspond to the top two ones in Fig. 6(f).
The original level-3 encoder shown in Fig. 8(a) is similar to the level-2 encoder in Fig. 6(a), but different implementations based on the Steane method [114] [Figs. 8(b) and 8(c)] are adopted for the error-detection gadgets [104]. The proposed encoder shown in Fig. 8(d) is similar to the level-2 one in Fig. 6(f), but it does not detect logical $X$ errors. This is enough if we use the level-3 logical qubits for quantum computation (not use for level-4 logical state preparation). Also, we perform simultaneous $Z$-error and $X$-error detection for not only level-1 blocks using the method in Fig. 6(g) or 6(h), but also a level-2 block using the method in Fig. 8(e) or 8(f).
@@
for the proposed encoder. (Note that the latter formula holds only for $n_{1},n_{2},n_{3}=4,6$}.) Thus, the proposed encoder can reduce the overhead by $2n_{2}n_{1}+7n_{3}+5n_{2}/2+3n_{1}+15-n_{3}n_{1}/2$. This is about 60% of the original overhead for all the eight level-3 MHC codes. Thus, the proposed encoder realizes about 60% reduction of overhead.
@@ V Logical CNOT performance
The logical-CNOT error probability is estimated as follows [104]. First, two error-free logical Bell pairs are prepared by performing error-free logical Hadamard and CNOT gates on error-free logical all-zero states. Next, we perform ten times transversal physical CNOT gates followed by error-correcting teleportation [16, 104, 116] on the first blocks of the two Bell pairs, where we use the encoders in Sec. IV and physical operations are faulty according to the above error model. Then, we disentangle the logical Bell pairs by performing error-free logical Hadamard and CNOT gates. Last, we measure all the physical qubits in the $Z$ basis and decode the results. We define the block error probability as $p_{\mathrm{block}}=1-(1-p_{10})^{1/10}$}, where $p_{10}$ is the probability that at least one of the logical-qubit values obtained by the decoding is 1. The error probability per logical CNOT, $p_{\mathrm{CNOT}}$, is also defined as $p_{\mathrm{CNOT}}=1-(1-p_{\mathrm{block}})^{1/K^{(3)}}$, where $K^{(3)}$ is the number of physical qubits in the level-3 block.
@@
Figure 9 shows the results for the four level-3 MHC codes with the original and proposed encoders. All the exponents obtained by the power-function fitting are close to d/2=4{d/2=4} (d=8{d=8} is the code distance). This suggests the fault tolerance of the original and proposed encoders. Also, all the solid lines are a little lower than the corresponding dashed lines. This means that the proposed encoders are a little better than the original ones. Most importantly, $D_{6,4,4}$ achieve the lowest values for both $p_{\mathrm{block}}$ and $p_{\mathrm{CNOT}}$. This concludes that $D_{6,4,4}$ is the best of the level-3 MHC codes.
@@
Figure 9: Numerical results of logical-CNOT error probability for level-3 MHC codes. (a–d) Log-log plots of the block error probability $p_{\mathrm{block}}$ for $D_{4,4,4}$, $D_{6,4,4}$, $D_{6,6,4}$, and $D_{6,6,6}$, respectively. $p_{\mathrm{block}}$ is the probability that at least one logical CNOT fails per block (see the main text for its definition). The circles and triangles represent the results with the proposed and original encoders, respectively. The solid and dashed lines are the fitting results with the power function $\beta p_{\mathrm{circ}}^{\alpha}$ ($\alpha$ and $\beta$ are fitting parameters) to the circles and triangles, respectively. In each panel, $p_{\mathrm{circ}}^{\alpha}$ shows the fitting result of $\alpha$. (e) $p_{\mathrm{block}}$ at $p_{\mathrm{circ}}=2\times 10^{-4}$}. (f) Logical-CNOT error probability $p_{\mathrm{CNOT}}$ at $p_{\mathrm{circ}}=2\times 10^{-4}$}. $p_{\mathrm{CNOT}}$ is defined as $p_{\mathrm{CNOT}}=1-(1-p_{\mathrm{block}})^{1/K^{(3)}}$, where $K^{(3)}$ is the number of physical qubits in the level-3 block.
@@
Figure 10 shows the total number of physical qubits necessary for preparing a logial all-zero state, the increase due to the repetition when errors are detected is taken into account. The results clearly show that the proposed encoders can reduces the overheads compared with the original encoders even when the error-detection effect is included.
@@
Figure 10: Total number of physical qubits necessary for preparing a logial all-zero state for level-3 MHC codes. (a–d) Semilog plots of the total number of physical qubits for $D_{4,4,4}$, $D_{6,4,4}$, $D_{6,6,4}$, and $D_{6,6,6}$, respectively. The solid and dashed curves show the results for the proposed and original encoders, respectively.

### v2

@@ Abstract
Many-hypercube codes [H. Goto, Sci. Adv. 10, eadp6388 (2024)], concatenated [[n,n2,2]]{[[n,n-2,2]]} quantum error-detecting codes ($n$ is even), have recently been proposed as high-rate quantum codes suitable for fault-tolerant quantum computing. While the original many-hypercube codes with n=6{n=6} can achieve remarkably high encoding rates (about 30% and 20% at concatenation levels 3 and 4, respectively), they have large code block sizes at high levels (216 and 1296 physical qubits per block at levels 3 and 4, respectively), making not only experimental realization difficult but also logical error rates per block high. Toward earlier experimental realization and lower logical error rates, here we comprehensively investigate smaller many-hypercube codes with $[[6,4,2]]$ and/or $[[4,2,2]]$ codes, where, e.g., $D_{6,4,4}$ denotes the many-hypercube code using $[[6,4,2]]$ at level 1 and $[[4,2,2]]$ at levels 2 and 3. As a result, we found a notable fact that $D_{6,4,4}$ ($D_{6,6,4,4}$) can achieve lower block error rates than $D_{4,4,4}$ ($D_{4,4,4,4}$), despite its higher encoding rate. Focusing on level 3, we also developed efficient fault-tolerant encoders realizing about 60% overhead reduction while maintaining or even improving the performance, compared to the original design. Using them, we numerically confirmed that $D_{6,4,4}$ also achieves the best performance for logical controlled-NOT gates in a circuit-level noise model. These results will be useful for early experimental realization of fault-tolerant quantum computing with high-rate quantum codes.
@@ I Introduction
Topological codes [24], in particular, the surface code [25, 26], have become more popular since it was shown that the surface code can achieve comparable thresholds only using local operations [27, 28, 29, 30, 31]. Moreover, a compact version called the rotated surface code and a simple two-qubit gate approach called lattice surgery were also proposed [32]. The surface-code approach to FTQC is particularly suitable for superconducting-circuit implementations, leading to recent experimental demonstrations of the surface code with superconducting circuits [33, 34, 35, 36].
@@
The (rotated) surface code encodes a single logical qubit into $d^{2}$ physical qubits, where $d$ is the code distance. Thus, the encoding rate $R=1/d^{2}$ becomes low for large code block sizes, resulting in large resource overheads. To address this overhead issue, quantum low-density parity-check (qLDPC) codes [43, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], which can achieve a constant encoding rate even for large code distances, have recently attracted much attention [48, 49, 50, 51, 52, 53]. This expectation is enhanced by recent experimental advances, in particular, in ion-trap [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67] and neutral-atom [68, 69, 70, 71, 72, 73, 74, 75, 76, 77] platforms, because long-range connectivity beyond nearest-neighbor one necessary for high-rate codes such as qLDPC codes [78, 79] is experimentally feasible by moving qubits in these platforms. However, parallel execution of addressable logical gates with low overheads is challenging for qLDPC codes, the study on which is ongoing [80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102].
@@
Alternative high-rate codes, namely, non-LDPC high-rate quantum codes have also been proposed toward low-overhead FTQC [103, 104, 105, 106, 107, 108, 109, 110]. Among them, concatenated high-rate quantum codes are particularly promising in the sense that fault-tolerant logical-state preparation can be achieved by level-by-level encoding [16, 22, 64, 103, 104, 105]. In this work, we focus on concatenated [[n,n2,2]]{[[n,n-2,2]]} quantum error-detecting codes ($n$ is even), which are called many-hypercube (MHC) codes because their structures can be described by using many $L$-dimensional hypercubes ($L$ is the concatenation level) corresponding to logical qubits (logical Pauli operators) [104, 107, 109]. Note that the error-detecting codes, which are also called iceberg codes [60], are the simplest, highest-rate Calderbank-Shor-Steane (CSS) codes [3, 4, 5]. In the original work [104], $[[6,4,2]]$ was adopted as the base code because of its good balance between the encoding rate and the code block size. Recently, smaller MHC codes with $[[4,2,2]]$ have been studied [107, 109]. Remarkably, the smallest level-2 MHC code $4^{2},2^{2},2^{2}]]$ has been realized experimentally using neutral atoms [77]. Smaller block sizes will lead to not only earlier experimental realization, but also lower logical error rates per block, though the encoding rates become lower. For this reason, in this work we investigate the MHC codes with $[[6,4,2]]$ and/or $[[4,2,2]]$, where the MHC code using $n_{L},n_{L}-2,2]]$} at level $L$ is denoted by $D_{n_{1},n_{2},\ldots}$ [107]. As a result, we found that $D_{6,4,4}$ ($D_{6,6,4,4}$) can achieve lower block error rates than $D_{4,4,4}$ ($D_{4,4,4,4}$), despite its higher encoding rate. In addition, it turned out that $D_{4,4,6,6}$ is the worst at level 4. This is a notable result, because in a previous study [107], it was assumed that we should use smaller codes at lower levels, and consequently $D_{4,4,6,6}$ was chosen in the previous study. We also develop efficient fault-tolerant encoders for level-3 MHC codes, which can reduce overheads by about 60% compared to the original design [104]. Using them, we confirmed that $D_{6,4,4}$ can also achieve the highest performance for logical CNOT gates at level 3.
@@
This paper is organized as follows. In Sec. II, we explain the definitions of the MHC codes in detail using figures visualizing their structures. In Sec. III, we show numerical results for block error rates using a bit-flip (depolarizing) error model. In Sec. IV, we present our new encoders for level-3 MHC codes comparing them with the original ones. In Sec. V, we show numerical results for logical-CNOT performance at level 3 using the new and original encoders. Finally, we summarize the results and discuss the outlook in Sec. VI.
@@ II.1 $D_{4}$ and $D_{6}$
Similarly to the case of $D_{4}$, these can be visualized as shown in Fig. 1(b). There are different definitions of [[n,n2,2]][[n,n-2,2]] codes [60], but we adopt the above definitions because of the simple visualization [104].
@@ II.4 Level-4 MHC codes
The level-4 MHC codes $D_{n_{1},n_{2},n_{3},4}$ and $D_{n_{1},n_{2},n_{3},6}$ are defined by similar equations to those for $D_{n_{1},n_{2},4}$ and $D_{n_{1},n_{2},6}$, respectively, in the last section. By introducing the fourth virtual axis denoted by $t$ and stacking level-3 $D_{n_{1},n_{2},n_{3}}$ blocks along the $t$ axis, each logical Pauli operator of the level-4 MHC codes corresponds to a four-dimensional hypercube, where $\mathrm{Q}_{i,j,k}$ of the $l$th $D_{n_{1},n_{2},n_{3}}$ block is placed at $(x,y,z,t)=(i,j,k,l)$. This is the origin of the name of MHC codes.
@@ III Performance comparison of MHC codes
Thus, the initial state can be prepared using the encoders in Figs. 5(a) and 5(b), respectively, where we assume no errors on all the operations. Next, we induce bit-flip errors on the physical qubits, then measure all the physical qubits in the $Z$ basis, and finally decode the measurement results using the level-by-level minimum-distance decoder developed in the original work [104, 111]. Unless all the logical-qubit values obtained by the decoding are 0, we consider that the decoding fails, leading to the block error probability. In this work, we used a Python package called Stim [115] for quantum circuit simulations.
@@
We also evaluated the performance of level-3 and level-4 MHC codes at $p_{\mathrm{flip}}=1\%$} and 2%, respectively, the results of which are shown in Figs. 5(b) and 5(c), respectively. As in the case of level 2, our results clearly suggest that we should use larger-size codes for lower levels for MHC codes, contrary to the previous assumption [107]. More importantly, Figs. 5(b) and 5(c) provide another counterintuitive result that the highest-performance MHC code is not the smallest-size one. The best MHC code is $D_{6,4,4}$ at level 3 and $D_{6,6,4,4}$ at level 4. (The level-4 MHC code chosen in a previous work [107], $D_{4,4,6,6}$, is the worst.) This is the main result of this work.
@@ IV Proposed encoders of level-3 MHC codes
$=N^{(2)}_{n_{1},n_{2}}+n_{2}+3n_{1}/2-2$ (4)
for the proposed encoder. (Note that the latter formula holds only for $n_{1},n_{2}=4,6$}.) Here, $N^{(2)}_{n_{1},n_{2}}=n_{2}n_{1}$ on the right-hand side is the number of physical qubits in the level-2 code block, and the other terms are the number of ancilla qubits, namely, overheads. Thus, the proposed encoder can reduce the overhead by $7-n_{1}/2$.
@@
Figure 7: Direct measurements of level-2 logical $Z$ operators. Here we show only the cases of $D_{n,4}$ (n=4,6{n=4,6}), but the cases of $D_{n,6}$ are similar. The small squares are physical ancilla qubits used for the measurements, which are initially set in $|0\rangle$. (a,b) Gate sequence for the logical $Z$ measurements used in the $D_{4,4}$ (a) and $D_{6,4}$ (b) encoders. Note that the bottom two level-1 blocks correspond to the top two ones in Fig. 6(f).
The original level-3 encoder shown in Fig. 8(a) is similar to the level-2 one in Fig. 6(a), but different implementations based on the Steane method [114] [Figs. 8(b) and 8(c)] are adopted for the error-detection gadgets [104]. The proposed encoder shown in Fig. 8(d) is similar to the level-2 one in Fig. 6(f), but it does not detect logical $X$ errors. This is enough if we use the level-3 logical qubits for quantum computation (not use for level-4 logical state preparation). Also, we perform simultaneous $Z$-error and $X$-error detection for not only level-1 blocks using the method in Fig. 6(g) or 6(h), but also a level-2 block using the method in Fig. 8(e) or 8(f).
@@
for the proposed encoder, where $N^{(3)}_{n_{1},n_{2},n_{3}}=n_{1}n_{2}n_{3}$. (Note that the latter formula holds only for $n_{1},n_{2},n_{3}=4,6$.) Thus, the proposed encoder can reduce the overhead by $2n_{2}n_{1}+7n_{3}+5n_{2}/2+3n_{1}+15-n_{3}n_{1}/2$. This is about 60% of the original overhead for all the eight level-3 MHC codes. Thus, the proposed encoder realizes about 60% reduction of overhead.
@@ V Logical CNOT performance
The logical-CNOT error probability is estimated as follows [104]. First, two error-free logical Bell pairs are prepared by performing error-free logical Hadamard and CNOT gates on error-free logical all-zero states. Next, we perform ten times transversal physical CNOT gates followed by error-correcting teleportation [16, 104, 116] on the first blocks of the two Bell-pair blocks, where we use the encoders in Sec. IV and physical operations are faulty according to the above error model. Then, we disentangle the logical Bell pairs by performing error-free logical Hadamard and CNOT gates. Last, we measure all the physical qubits in the $Z$ basis and decode the results. We define the block error probability as $p_{\mathrm{block}}=1-(1-p_{10})^{1/10}$}, where $p_{10}$ is the probability that at least one of the logical-qubit values obtained by the decoding is 1. The error probability per logical CNOT, $p_{\mathrm{CNOT}}$, is also defined as $p_{\mathrm{CNOT}}=1-(1-p_{\mathrm{block}})^{1/K^{(3)}}$, where $K^{(3)}$ is the number of logical qubits in the level-3 block.
@@
Figure 9 shows the results for the four level-3 MHC codes with the original and proposed encoders. All the exponents obtained by the power-function fitting are close to d/2=4{d/2=4} (d=8{d=8} is the code distance). This suggests the fault tolerance of the original and proposed encoders. Also, all the solid lines are a little lower than the corresponding dashed lines. This means that the proposed encoders are a little better than the original ones. Most importantly, $D_{6,4,4}$ achieves the lowest values for both $p_{\mathrm{block}}$ and $p_{\mathrm{CNOT}}$. This concludes that $D_{6,4,4}$ is the best of the level-3 MHC codes.
@@
Figure 9: Numerical results of logical-CNOT error probability for level-3 MHC codes. (a–d) Log-log plots of the block error probability $p_{\mathrm{block}}$ for $D_{4,4,4}$, $D_{6,4,4}$, $D_{6,6,4}$, and $D_{6,6,6}$, respectively. $p_{\mathrm{block}}$ is the probability that at least one logical CNOT fails per block (see the main text for its definition). The circles and triangles represent the results with the proposed and original encoders, respectively. The solid and dashed lines are the fitting results with the power function $\beta p_{\mathrm{circ}}^{\alpha}$ ($\alpha$ and $\beta$ are fitting parameters) to the circles and triangles, respectively. In each panel, $p_{\mathrm{circ}}^{\alpha}$ shows the fitting result of $\alpha$. (e) $p_{\mathrm{block}}$ at $p_{\mathrm{circ}}=2\times 10^{-4}$}. (f) Logical-CNOT error probability $p_{\mathrm{CNOT}}$ at $p_{\mathrm{circ}}=2\times 10^{-4}$}. $p_{\mathrm{CNOT}}$ is defined as $p_{\mathrm{CNOT}}=1-(1-p_{\mathrm{block}})^{1/K^{(3)}}$, where $K^{(3)}$ is the number of logical qubits in the level-3 block.
@@
Figure 10 shows the total number of physical qubits necessary for preparing a logical all-zero state, taking the repetition due to error detection into account. The results clearly show that the proposed encoders can reduces the overheads compared with the original encoders even when the error-detection effect is included.
@@
Figure 10: Total number of physical qubits necessary for preparing a logical all-zero state for level-3 MHC codes. (a–d) Semi-log plots of the total number of physical qubits for $D_{4,4,4}$, $D_{6,4,4}$, $D_{6,6,4}$, and $D_{6,6,6}$, respectively. The solid and dashed curves show the results for the proposed and original encoders, respectively.